### v1

@@ Abstract
Backbone-assisted assembly processes – such as protein folding – allow the assembly of a large number of structures with high accuracy from only a small handful of fundamental building blocks. We aim to explore general principles underlying this phenomenon by studying variants of the temperature-1 abstract tile assembly model (aTAM). We consider the existence of finite sets of tile types that can deterministically assemble any shape producible by a given assembly model; we call such tile type sets universal assembly kits. Our first model, which we call the “backboned aTAM”, generates backbone-assisted assembly by forcing tiles to be be added to lattice positions neighbouring the immediately preceding tile, using a predetermined sequence of tile types. We demonstrate the existence of universal assembly kit for the backboned aTAM, and show that the existence of this set is maintained even under stringent restrictions to the rules of assembly. We hypothesise that the advantage of the backboned aTAM relative to the conventional aTAM – which does not possess such a set – is in part due to the specification of a sequence, and in part due to the geometric restrictions imposed by the backbone. To explore this intuition, we develop a second model call the “sequenced aTAM”. The sequenced aTAM uses a predetermined sequence of tiles, but does not constrain a tile to neighbour the immediately preceding tiles. We prove that this model has no universal assembly kit. The lack of such a kit is surprising, given that the number of tile sequences of length $N$ scales faster than the number of producible shapes of size $N$ for a sufficiently large – but finite – set of tiles.
keywords:
self-assembly, abstract Tile assembly model, macromolecular folding
@@ 1.1 Motivation and Aims
Molecular self-assembly, the process through which molecules adopt defined configurations in the absence of external control, is the foundation of biochemical complexity. The most basic form of self-assembly involves multiple subunits in solution spontaneously coming together to create a more complex structure, with that structure exclusively determined by the interactions between the subunits. The archetypal example is the joining of protein subunits to produce a larger quarternary structure ([12]). Backbone-assisted assembly is a distinct process through which biology constructs complex structures. Here, the subunits are first assembled into a sequence connected by a backbone before folding into the final structure. Crucially, the sequence of subunits is not determined by the interactions between the subunits themselves. A quintessential example is the folding of polypeptide sequences into protein secondary and tertiary structures, with the sequence specified by an mRNA template ([2]).
Backbone-assisted assembly allows control over which molecular building blocks are incorporated into a structure, and where. The consequences are evident in protein folding. Here, a large number of folded polypeptide structures are reliably selected from a vast space of possibilities, while utilizing only 20 amino acid subunits. If these same amino acids were instead assembled without first being polymerized into a sequence, the resultant structures would be non-specific ([26]). Conversely, if we did have a set of 20 building blocks that spontaneously formed a specific structure, then they would struggle to form alternative structures containing the same 20 basic subunits. In general, the spontaneous assembly of a large number of specific structures necessitates a number of building blocks that grows with the number of targets if specificity is to be maintained ([24]).
With the advent of DNA nanotechnology, the last 20 years have seen the rise of synthetic systems capable of assembling into complicated structures ([23, 27]). Although DNA itself is a copolymer of backbone-linked nucleotides, the key features of backbone-facilitated assembly are typically not explored in these systems. Many rely on the free assembly of short DNA oligonucleotide strands ([30, 14, 22]). The technique of DNA origami ([23]) involves a long scaffold strand that is folded by short staples that bind to non-contiguous domains. These domains act like subunits embedded in a backbone chain, with their interactions mediated by the staples. However, with few exceptions ([32, 9]) the staples are intended to connect a unique pair of domains. Moreover, re-ordered domains are not used to drive the formation of distinct, well-defined structures. The question of how a small number of basic units can direct the assembly of many distinct structures does not, therefore, arise.
Backbone-assisted assembly schemes could help to reduce assembly complexity in theses synthetic systems by encoding a greater portion of the target structure through the backbone, reducing the number of subunits that must be fabricated. Progress in this space include developments in single-stranded nucleic acid nanotechnology ([28, 11, 33, 16, 13]) as well as alternative techniques such as programmable droplets ([19]). While promising, the sophistication of shapes accessible to existing systems is generally lower than those designed with conventional methods. A theory of backbone-assisted assembly could potentially aid in the development of more powerful artificial backbone-assisted assembly schemes.
While models of the backbone-assisted assembly exist ([17],[10]), previous work hasn’t focused explicitly on the backbone’s importance relative to systems with no backbone, and existing models do not have straightforward non-backboned analogs for direct comparison. In this work, we develop a model based on the temperature-1 abstract Tile Assembly Model (aTAM), to build towards an understanding of general principles that give backbone-assisted assembly systems their unique advantages.
1.2 Review of aTAM
The aTAM is a model of computation via self-assembly ([24]). The fundamental building blocks of the aTAM are tiles, with each face associated with a numbered glue. An aTAM instance is a tuple $(T,g,A_{0},\tau)$, where $T$ is a set of tile types, $g$ is a strength function that maps tuples of glue types onto integer values representing the strengths of their interaction. $A_{0}$ is a starting seed configuration (a set of tiles bound together in space) and $\tau$ is the temperature. Assembly occurs in discrete time steps, such that at each step, a tile is stuck onto the growing configuration. A tile can only be stuck onto the growing configuration if the sum of strengths of interactions formed between its glues and glues of the growing configuration meets or exceeds the temperature of the aTAM system.
For the purpose of studying assembly, it is natural to consider the number of tile types necessary to construct a particular shape deterministically, that is, without any chance of constructing off-target shapes. Existing work (largely on temperature-2 aTAM) generally considers scaled versions of shapes as equivalent ([29]), and the notion of a scaling factor, the scale at which some tile type set assembles a shape, arises as a result. By allowing shape scaling, the related concept of intrinsic universality arises. Intrinsic universality refers to the ability of an assembly scheme to not only assemble any shape, but simulate any other instance of the assembly scheme, in the sense of being able to replicate sequences of tile addition at some scale factor ([31]). The temperature-2 is aTAM is known to be intrinsically universal ([7, 6, 31]), whereas the temperature-1 aTAM is not intrinsically universal ([8, 21, 20]).
Intrinsic universality is a measure of the computational power of an assembly scheme. We have opted to use a different, but related, measure of the power of assembly schemes. We will consider whether particular assembly schemes admit a finite set of tile types that enable the deterministic assembly of all shapes achievable (achievable here means that there exists some input to the assembly system that constructs the given shape) by the given assembly scheme, and we call such a tile type set a universal assembly kit. In contrast to intrinsic universality, where non-trivial starting seeds are allowed, the definition of a universal assembly kit also explicitly disallows the use of any non-trivial starting seed. Further, we will be relying on assembly schemes which explicitly allow the rotation of tiles. These changes are consistent with our aims, as we are concerned with macromolecular assembly rather than using assembly for computation.
The existence of a universal assembly kit is weaker than intrinsic universality in the sense that such a set may not be able to simulate exact sequences of tile addition, but stronger in that shapes must be constructed at their exact scale. Since our models were developed from temperature-1 aTAM, we will be comparing results to the temperature-1 aTAM as a base. The non-existence of a universal assembly kit for the temperature-1 aTAM is provable via pumpability arguments on rectangles ([24]). Algorithms do exist, however, for the minimization of tile type sets to deterministically construct particular shapes using variants of temperature-1 aTAM, via graph theoretic ([1]) and boolean satisfiability ([25, 4]) approaches.
@@ 1.3 Summary of Results
We develop a model based on the temperature-1 aTAM that simulates a backbone-assisted assembly process, which we call the backboned aTAM. We consider a variant of the aTAM that allows rotations ([6]) and for generality we allow mismatches to have negative strength, leading to mismatches that may block the addition of tiles (as in ([18])). Like the base aTAM, the fundamental building blocks of our system are tiles with labelled faces. However, tiles can only be added in a specified order of tile types, and added tiles must neighbour the last added tile. The model can be thought of as mimicking idealized co-transcriptional folding; the added sequence information present in the backboned aTAM means it can be more powerful at selectively producing specific shapes, allowing for a more repetitive use of tiles.
We provide here an informal definition of a universal assembly kit as this is vital to summarizing our results (a more thorough definition is presented in Section 3.1). Let an assembly scheme be a set of rules that constructs shapes from tiles with numbered faces and possible additional inputs (in our case, these additional inputs are usually a sequence in which tiles can be added). Then, a universal assembly kit is a finite set of tile types and interaction rules that allow any shape to be constructed deterministically, if that shape can be constructed at all by the assembly scheme (i.e. without imposing constraints on the set of tile types that may be used, some input to the assembly scheme exists that constructs the shape).
Our first main result is the existence of a universal assembly kit for the backboned aTAM.
@@
The construction we arrive at to prove Theorem 3.2 is in some sense artificial as it relies exclusively on backbone routing to define shapes while disregarding non-backbone interactions. To circumvent this problem, and get closer to protein folding in which interactions with non-neighbouring amino acids are essential in determining the fold, we show that the existence of a universal assembly kit is preserved even when all adjacent tile faces are required to have attractive glue interactions. This finding recalls existing work on the aTAM, where mismatches and rotations were shown to have a weak effect on computational power ([18]).
@@
A universal assembly kit for the backboned aTAM exists such that the strength function $g$ associated with the assembly kit does not encode any neutral interactions.
These two results contrast with the temperature-1 base aTAM, where a universal assembly kit does not exist ([24]). To help us disentangle the effects of specifying a sequence and restricting the geometry of tile placement via a backbone, we develop a model we call the sequenced aTAM, where tiles are added in a predetermined sequence, as in the backboned aTAM, but added tiles are no longer constrained to neighbor preceding tiles. While not reflective of any autonomous real world process, this construct serves as a null model for comparison to the backboned aTAM. We find that unlike the backboned aTAM, a universal assembly kit does not exist in the case of the sequenced aTAM.
@@
This result is surprising given the known exponential scaling of the set of aTAM-producible shapes of a given size, since it is always possible to chose a finite tile set with a faster exponential scaling of the number of sequences of a given length ([15]). Although this loss of a universal assembly kit is due to the loss of the geometric constraints imposed by the backbone, we cannot conclude that the presence of the backbone (in addition to the sequence) makes the backboned aTAM more powerful than the sequenced aTAM. This complication arises because the backboned aTAM has a smaller space of achievable shapes than the sequenced aTAM. Indeed, the sequenced aTAM can deterministically assemble any shape that can be assembled by the backboned aTAM using a finite assembly kit (such a kit would be non-universal in the context of the sequenced aTAM). However, we have been unable to prove whether the sequenced aTAM can – like the backboned aTAM – do so even when we impose no neutral interactions. We conclude by presenting some partial results in this direction.
@@ 2 Model Definitions
We begin with definitions borrowed from the aTAM literature ([24, 8]). We consider a variant of aTAM where tiles are allowed to rotate ([6]). As tiles are allowed to rotate, we differentiate between an oriented tile, a 4-tuple of glue types $(\sigma_{N},\sigma_{E},\sigma_{S},\sigma_{W})$, and the orientation-free tile type which is an equivalence class of all cyclic permutations of any of its given tiles. In practice, the tile type can be associated with a default orientation, and the tile can be conceived of a tile type $\Theta$ placed in some specific orientation $\rho=\{N,E,S,W\}$ corresponding to the direction faced by the face that would face north in the default orientation.
@@
Figure 1: A diagram illustrating aTAM assembly. The fundamental building blocks of assembly are square tiles with numbered faces. At each step, a tile drawing from the assigned set of tile types (bottom) is added to a random position neighbouring an existing tile, with the possible locations being restricted to those where the resulting sum of strengths of interactions formed is greater than or equal 1 (for temperature-1 aTAM). Here, the glue pair $(1,2)$ are predetermined to have strength of 1. Note that we use a variant of aTAM where tiles can be rotated. Example assembly steps steps starting from the state at the top left of the figure are given in top centre and top right.
@@
A configuration is a partial function $A:\mathbb{Z}^{2}\rightarrow\mathbb{T}$, where $\mathbb{T}$ is the set of all possible tiles. $dom(A)$ is the set of points in configuration $A$ with a tile. A coordinate $z\notin dom(A)$ is empty in $A$. $A$ is a subconfiguration of $A^{\prime}$ if $dom(A)\subseteq dom(A^{\prime})$. For convenience, we frequently use single-tile configurations $a=(\Theta,\rho,z)$ for a tile type $\Theta$, an orientation $\rho$ and a coordinate $z$, which we call coordinated tiles. The empty configuration is defined as $A_{empty}$ such that $dom(A_{empty})=\emptyset$. The addition of configurations $A^{\prime\prime}=A+A^{\prime}$ is well-defined if $dom(A)\bigcap dom(A^{\prime})=\emptyset$, otherwise $A^{\prime\prime}=\infty$ ([24]). In the former case,
$A^{\prime\prime}(z)=\begin{cases}A(z)\;\>if\;\>z\in dom(A),\\ A^{\prime}(z)\;\>if\;\>z\in dom(A^{\prime}).\\ \end{cases}$
@@
The strength function is a partial function $g:\mathbb{Z}_{+}^{2}\rightarrow\mathbb{Z}$. The strength function determines the type of interaction between two glues. Two glues $\sigma$ and $\sigma^{\prime}$ have an attractive interaction if $g(\sigma,\sigma^{\prime})>0$, do not interact if $g(\sigma,\sigma^{\prime})=0$ (or have a neutral interaction) and have a repulsive interaction if $g(\sigma,\sigma^{\prime})<0$. Newly added tiles are allowed to form neutral or repulsive interactions, but only if the sum of interactions of each of their edges add up to greater than the temperature $\tau$, which we have set equal to $1$ for all the models that we consider.
@@
The aim of any instance of an aTAM system is to produce shapes. For the base aTAM, shapes are assembled through the addition of tiles drawing from the tile type set $T$ until no further tiles can be added (Refer to Figure 1 for an illustration, and reference ([24]) for formal definitions). We describe as follows the way in which our two models, the backboned aTAM and the sequenced aTAM assemble their shapes.
@@
We refer to the set of non-empty points $dom(A)$ as the oriented shape of a configuration $A$. We consider all outputs that can be transformed through rotations and translations as equivalent, hence we rely on a more general notion than oriented shape. A shape is thus an equivalence class of oriented shapes, containing oriented shapes that can be transformed to each other via rotations and translations, and the shape of a configuration is the shape to which its oriented shape belongs. We denote by $\simeq$ the equivalence relation that defines a shape, such that if two oriented shapes $S$ and $S^{\prime}$ have the same shape, then $S\simeq S^{\prime}$. Hence, the only ambiguity in shape is overall rotation or translation, which do not violate the equivalence class.
@@
Intuitively, the backboned aTAM aims to mimic a cotranslational folding system. Unlike the base aTAM, tiles can only be added in a predetermined order (a sequence of tile types is provided as an input to the assembly system), and any added tiles must neighbour the last added tile (Figure 2). Formally, a backboned aTAM instance is a 3-tuple $(\mathcal{A},Q,g)$ where $\mathcal{A}$ is an initial configuration, $Q$ is a sequence of tile types and $g$ is a strength function.
@@
A specific instance of the backboned aTAM can generate trajectories through the addition of tiles with types given by the sequence $Q$ of the backboned aTAM system, while obeying the strength function $g$ and rules about tile placement. A trajectory that can be generated by a backboned aTAM system is said to be complete with respect to the backboned aTAM system. We can now formally define a complete trajectory.
@@
Consider a backboned aTAM instance $(\mathcal{A},Q,g)$, with the sequence of tile types $Q=(\Theta_{1},\Theta_{2},..,\Theta_{\omega})$. A trajectory $\Psi=(A_{0},A_{1},A_{2},..)$ is said to be complete with respect to $(\mathcal{A},Q,g)$ if the following hold.
  1. 1.
    Starting configurations are consistent, that is $A_{0}=\mathcal{A}$.
  2. 2.
    $A_{t}=A_{t-1}+a_{t}$ with $a_{t}$ chosen randomly under the constraints:
    1. (a)
      $A_{t}\neq\infty$ for any $t$
    2. (b)
    3. (c)
      Any coordinated tile in $A_{t}$ is such that the sum of strengths $\sum_{i}g(\sigma_{i},\sigma_{i}^{\prime})$, where $\sigma_{i}$ is the face $i$ of the tile, $\sigma^{\prime}_{i}$ is the tile face adjacent to $\sigma_{i}$ and $g(\sigma_{1},\sigma_{1}^{\prime})=0$ if $\sigma_{i}^{\prime}$ belongs to an empty coordinated tile (A configuration $A_{t}$ that obeys this assumption for some interaction function $g$ is called g-valid).
    4. (d)
  3. 3.
    The trajectory terminates upon reaching $A_{\omega}$ or when no such single tile configuration $a_{t}$ can be added. In the latter case, the trajectory is said to have been prematurely terminated.
By this definition, the first tile added in an empty configuration can be placed in any coordinate, at any orientation. The set of assembled configurations for an instance of the backboned aTAM is defined as the set containing all terminal configurations for all complete trajectories generated by the backboned aTAM instance. The set of assembled shapes for a backboned aTAM instance is thus the set containing all shapes of assembled configurations of the backboned aTAM instance.
We further define a model that we call the sequenced aTAM, which serves as a null model against which features of the backboned aTAM can be compared. The definition of a sequenced aTAM instance, along with the definition of its complete trajectories, are as follows. A sequenced aTAM instance is a 3-tuple $(\mathcal{A},Q,g)$ (and hence it has identical descriptors to a backboned aTAM instance). The key difference lies in the definition of a complete trajectory for the sequenced aTAM, which does not mandate that tiles are placed adjacent to the previous tile.
Definition 3.
Consider a sequenced aTAM instance $(\mathcal{A},Q,g)$, with the sequence of tile types $Q=(\Theta_{1},\Theta_{2},..,\Theta_{\omega})$. A trajectory $\Psi=(A_{0},A_{1},A_{2},..)$ is said to be complete with respect to $(\mathcal{A},Q,g)$ if the following hold.
@@
      Any coordinated tile in $A_{t}$ is such that the sum of strengths $\sum_{i}g(\sigma_{i},\sigma_{i}^{\prime})$, where $\sigma_{i}$ is the face $i$ of the tile, $\sigma^{\prime}_{i}$ is the tile face adjacent to $\sigma_{i}$ and $g(\sigma_{1},\sigma_{1}^{\prime})=0$ if $\sigma_{i}^{\prime}$ belongs to an empty coordinated tile (A configuration $A_{t}$ that obeys this assumption for some interaction function $g$ is called g-valid).

### v2

@@ Abstract
Sequence-directed assembly processes – such as protein folding – allow the assembly of a large number of structures with high accuracy from only a small handful of fundamental building blocks. We aim to explore how efficiently sequence information can be used to direct assembly by studying variants of the temperature-1 abstract tile assembly model (aTAM). We ask whether, for each variant, their exists a finite set of tile types that can deterministically assemble any shape producible by a given assembly model; we call such tile type sets “universal assembly kits”. Our first model, which we call the “backboned aTAM”, generates backbone-assisted assembly by forcing tiles to be added to lattice positions neighbouring the immediately preceding tile, using a predetermined sequence of tile types. We demonstrate the existence of universal assembly kit for the backboned aTAM, and show that the existence of this set is maintained even under stringent restrictions to the rules of assembly. We compare these results to a less constrained model that we call sequenced aTAM, which also uses a predetermined sequence of tiles, but does not constrain a tile to neighbour the immediately preceding tiles. We prove that this model has no universal assembly kit in the stringent case. The lack of such a kit is surprising, given that the number of tile sequences of length $N$ scales faster than both the number and worst-case Kolmogorov complexity of producible shapes of size $N$ for a sufficiently large – but finite – set of tiles. Our results demonstrate the importance of physical mechanisms, and specifically geometric constraints, in facilitating efficient use of the information in molecular programs for structure assembly.
@@ 1.1 Motivation and Aims
Biological systems apply a wide range of molecular assembly maps that accept some generalised “genotype” as input and produce some phenotpye as output – for example, the RNA genotype-phenotype map (?) – to achieve their biochemical complexity. One type of assembly map involves multiple subunits in solution spontaneously coming together to create a more complex structure, with that structure exclusively determined by the interactions between the subunits; here, the input genotype is an unordered set of building blocks. The archetypal example is the joining of protein subunits to produce a larger quaternary structure (?). Contrast this map type with sequence-directed assembly processes where subunits are first assembled into a sequence connected by a backbone, before folding into the final structure; here, the input genotype is a sequence of building blocs. A quintessential example is the folding of polypeptide sequences into protein secondary and tertiary structures, with the sequence specified by an mRNA template (?).
Sequence-directed assembly comes with some unique advantages. For example, it tends to utilize a relatively small set of building blocks to create a vast array of specific structures through sequence variation. This advantage is evident in protein folding, where just 20 amino acid types can produce countless distinct functional structures (???). Conversely, non-sequenced assembly typically requires a larger set of building blocks to achieve similar structural specificity (?). If we had a set of 20 building blocks that spontaneously formed a specific structure, they would struggle to form alternative structures containing the same basic subunits (?). While this argument is intuitive, exactly which features facilitate the efficient use of building blocks for a given sequence-directed assembly map is less well understood.
The importance of sequence-directed assembly in biology is appreciated, and a number of models have been put forward to investigate different aspects of sequence-directed assembly. Early examples include 1D lattice models, such as the Lifson-Roig model (?), which were used to characterize helix-coild secondary structure phase transitions. The HP-Lattice model (?), an approximate model of protein folding, has been used to understand the role of hydrophobic and hydrophilic amino acids in protein folding. RNA secondary structure folding has been used to evaluate the genotype-phenotype map properties of sequence-directed assembly maps due to the speed with which low free-energy configurations can be identified (?). The computational capabilities of backbone-directed assembly have been explored using the Oritatami model (???).
One important motivation for investigating these theoretical models of folding is to further develop artificial molecular assembly systems, including DNA nanotechnology-based approaches (??). Notably, while DNA itself is a copolymer of backbone-linked nucleotides, DNA assembly systems utilize both sequence-directed and sequence-free assembly. Some assembly systems rely primarily on the free assembly of short oligonucleotide strands (???). DNA origami (?) represents an interesting hybrid approach, using a long scaffold strand folded by short staples that bind to non-contiguous domains; systems with reusable or reconfigurable modules attempt to more heavily exploit features of sequence-directed assembly (??). Single-stranded nucleic acid nanotechnology (?????) and alternative techniques such as programmable droplets (?) take greater advantage of sequence-directed assembly.
The models we explore in this work are based on the polyomino assembly model applied in (?), itself based on the temperature-1 abstract Tile Assembly Model (aTAM) (?) allowing for tile rotations (?). aTAM variants are frequently employed to explore the computational capabilities of self-assembly systems. In particular, much of the existing literature on the aTAM has focused on the question of intrinsic universality, the ability of an assembly system to simulate any other instance of the assembly scheme at some scaling factor (?). The temperature-1 aTAM is not intrinsically universal (???), while the temperature-2 aTAM is (???). Other work has investigated the minimum number of tiles required to produce specific shapes (???) from some starting seed and at some scale factor. Yet another line of literature considers minimal tilesets without starting seeds or scale factors (???); these works consider the physical properties of the resulting maps from the “genotype” of tile types to the assembly phenotype, and their relationship to the Kolmogorov complexities of polyominoes. This line of work has additionally resulted in Boolean Satisfiability algorithms for tile minimization (??).
In this paper, we compare different sequence-directed assembly models and consider their ability to deterministically assemble complex shapes using only a finite set of tiles; in doing so, we hope to develop an understanding of the features of efficient sequence-directed assembly maps. Our choice of the aTAM as a base model is motivated by this aim, as the underlying rules of the aTAM are simple enough that we can construct toy models that isolate specific features of real-world, sequence-directed assembly maps while selectively omitting others. For these toy models of assembly, we consider whether the information in a sequence can be used to direct the assembly of any possible shape, and whether this information can be used efficiently. We consider two measures of efficiency, one related to the scaling of the number of possible shapes with size, and the other related to the scaling in Kolmogorov complexity of shapes with size. Kolmogorov complexity is the length of a minimal program (under some universal language) required to produce a certain output (?). While the Kolmogorov complexity is formally incomputable, progress can be made by considering its upper bounds and scaling (???).
This paper is organized as follows. In Section 2, we present definitions of models and terms we use. In Section 3, we present our results (these are briefly summarized in Section 1.2) and discuss these results in light of the properties of our models. We discuss the conclusions and implications of this work in Section 4.
@@ 1.2 Summary of Results
We first consider an aTAM-based model that simulates a backbone-assisted assembly process, which we call the backboned aTAM. Like the base aTAM, the fundamental building blocks of our system are tiles with labeled faces. However, tiles can only be added in a specified order of tile types, and added tiles must neighbour the last added tile. The model can be thought of as mimicking idealized co-transcriptional folding; the added sequence information present in the backboned aTAM means it can be more powerful at selectively producing specific shapes, allowing for a more repetitive use of tiles.
We provide here an informal definition of a universal assembly kit as this concept is vital to summarizing our results (a more thorough definition is presented in Section 3.1). Let an assembly scheme be a set of rules that constructs shapes from tiles with numbered faces and possible additional inputs (in our case, these additional inputs are usually a sequence in which tiles can be added). Then, a universal assembly kit is a finite set of tile types and interaction rules that allow any finite shape to be constructed deterministically, if that shape can be constructed at all by the assembly scheme (i.e. without imposing constraints on the set of tile types that may be used, some input to the assembly scheme exists that constructs the shape).
Our first result is the existence of a universal assembly kit for the backboned aTAM.
@@
The construction we arrive at to prove Theorem 1 is in some sense artificial as it relies exclusively on backbone routing to define shapes while disregarding non-backbone interactions; interactions between tiles that are not connected by the backbone are all “neutral”. To circumvent this problem, and get closer to protein folding in which interactions with non-neighbouring amino acids are essential in determining the fold, we show that the existence of a universal assembly kit is preserved even when neutral interactions are forbidden and all adjacent tile faces are required to have attractive glue interactions. This finding recalls existing work on the aTAM, where mismatches and rotations were shown to have a weak effect on computational power (?).
@@
To help us understand the role of sequence information in very general cases, we develop a model that we call the sequenced aTAM, where tiles are added in a predetermined sequence, as in the backboned aTAM, but added tiles are no longer constrained to neighbor preceding tiles. While not reflective of an autonomous real-world process, such a model could describe a generalisation of the step-by-step synthesis protocols that are used to construct synthetic DNA and proteins (??). For our purposes, this construct serves as comparative model in which sequence information is not paired with geometric constraints. We find that unlike the backboned aTAM, a universal assembly kit does not exist in the case of the sequenced aTAM.
@@
These results on assembly kits can be interpreted in light of the scaling of shape space size and worst-case Kolmogorov complexity of the underlying shape space. Neither the shape space size nor the Kolmogorov complexity prohibit a universal assembly kit for either the backboned or sequenced aTAM. Indeed, the shape space provides a relatively small lower bound on the number of tile types. We find that, although the backboned aTAM does possess a universal assembly kit, requiring that all interactions are attractive makes the resultant “program” rather inefficient, due to both redundancy and failed assemblies. These inefficiencies are so large for the sequenced aTAM that no universal assembly kit is possible at all. This result hints at a role of the backbone in reducing assembly complexity beyond coupling shape assembly to a sequence.
Note, however, that the absence of backbone restrictions means that the sequenced aTAM can assemble shapes that are not formed by a single self-avoiding walk, whereas the backboned aTAM cannot, and therefore the sequenced aTAM has a larger structure space. We cannot therefore conclude that the backbone exclusively confers advantages to assembly. Moreover, the sequenced aTAM can, with a finite asssembly kit, deterministically assemble any shape in the structure space of the backboned aTAM, provided that neutral interactions are allowed. However, we have been unable to prove whether the sequenced aTAM can – like the backboned aTAM – do so even when we impose attractive inter-tile interactions. We conclude by presenting some partial results in this direction.
@@ 2 Model Definitions
We begin with definitions borrowed from the aTAM literature (??). As tiles are allowed to rotate, we differentiate between an oriented tile, which is a 4-tuple of glue types $(\sigma_{N},\sigma_{E},\sigma_{S},\sigma_{W})$, and the orientation-free tile type which is an equivalence class of all cyclic permutations of any of its given tiles. In practice, the tile type can be associated with a default orientation, and the tile can be conceived of a tile type $\Theta$ placed in some specific orientation $\rho=\{N,E,S,W\}$ corresponding to the direction faced by the face that would face north in the default orientation.
@@
Figure 1: A diagram illustrating aTAM assembly. The fundamental building blocks of assembly are square tiles with numbered faces. At each step, a tile drawing from the assigned set of tile types (bottom) is added to a random position neighbouring an existing tile, with the possible locations being restricted to those where the resulting sum of strengths of interactions formed is greater than or equal 1 (for temperature-1 aTAM). Here, the glue pair $(1,2)$ are predetermined to have strength of 1. Note that we use a variant of the aTAM in which tiles can be rotated. Example assembly steps steps starting from the state at the top left of the figure are given in top centre and top right.
@@
A configuration is a partial function $A:\mathbb{Z}^{2}\rightarrow\mathbb{T}$, where $\mathbb{T}$ is the set of all possible tiles. $dom(A)$ is the set of points in configuration $A$ with a tile. A coordinate $z\notin dom(A)$ is empty in $A$. $A$ is a subconfiguration of $A^{\prime}$ if $dom(A)\subseteq dom(A^{\prime})$. For convenience, we frequently use single-tile configurations $a=(\Theta,\rho,z)$ for a tile type $\Theta$, an orientation $\rho$ and a coordinate $z$, which we call coordinated tiles. The empty configuration is defined as $A_{empty}$ such that $dom(A_{empty})=\emptyset$. The addition of configurations $A^{\prime\prime}=A+A^{\prime}$ is well-defined if $dom(A)\bigcap dom(A^{\prime})=\emptyset$, otherwise $A^{\prime\prime}=\infty$ (?). In the former case,
$A^{\prime\prime}(z)=\cases{A}(z)\;\>if\;\>z\in dom(A),\\ A^{\prime}(z)\;\>if\;\>z\in dom(A^{\prime}).\$
@@
The strength function is a partial function $g:\mathbb{Z}_{+}^{2}\rightarrow\mathbb{Z}$. The strength function determines the type of interaction between two glues. Two glues $\sigma$ and $\sigma^{\prime}$ have an attractive interaction if $g(\sigma,\sigma^{\prime})>0$, do not interact if $g(\sigma,\sigma^{\prime})=0$ (also called a neutral interaction) and have a repulsive interaction if $g(\sigma,\sigma^{\prime})<0$. Newly added tiles are allowed to form neutral or repulsive interactions (see (?) for a further discussion of negative strengths), but only if the sum of interactions of each of their edges is equal to or greater than the temperature $\tau$, which we have set equal to $1$ for all the models that we consider.
@@
The aim of any instance of an aTAM system is to produce shapes. For the base aTAM, shapes are assembled through the addition of tiles drawing from the tile type set $T$ until no further tiles can be added (Refer to Figure 1 for an illustration, and reference (?) for formal definitions). We describe below the way in which our two models, the backboned aTAM and the sequenced aTAM, assemble their shapes.
@@
We refer to the set of non-empty points $dom(A)$ as the oriented shape of a configuration $A$. We consider all outputs that can be transformed through rotations and translations as equivalent; hence we rely on a more general notion than oriented shape. A shape is thus an equivalence class of oriented shapes, containing oriented shapes that can be transformed to each other via rotations and translations, and the shape of a configuration is the shape to which its oriented shape belongs. We denote by $\simeq$ the equivalence relation that defines a shape, such that if two oriented shapes $S$ and $S^{\prime}$ have the same shape, then $S\simeq S^{\prime}$. Hence, the only ambiguity in shape is overall rotation or translation, which do not violate the equivalence class.
@@
Intuitively, the backboned aTAM aims to mimic a cotranslational folding system. Unlike the base aTAM, tiles can only be added in a predetermined order (a sequence of tile types is provided as an input to the assembly system), and any added tiles must neighbour the last added tile (Figure 2). Formally, a backboned aTAM instance is a 3-tuple $(\mathcal{A},Q,g)$, where $\mathcal{A}$ is an initial configuration, $Q$ is an input sequence of tile types and $g$ is a strength function, that generates a set of complete trajectories, with trajectories being defined in Definition 1, and with the set of complete trajectories obeying Definition 2. Note that in our treatment of the interaction function $g$, we have allowed negative (respulsive) interactions. However, unlike the approach taken by (?) and (?), we do not allow negative interactions to displace existing tiles - rather, tiles that would form strong enough negative interactions to destabilize the configuration are simply blocked.
@@
A specific instance of the backboned aTAM can generate trajectories through the addition of tiles with types given by the sequence $Q$ of the backboned aTAM system, while obeying the strength function $g$ and rules about tile placement. A trajectory that can be generated by a backboned aTAM system is said to be complete with respect to the backboned aTAM system. We can now formally define a complete trajectory. Note that in using cuts and cut-edges, we follow the approach in (?) and (?).
@@
Consider a backboned aTAM instance $(\mathcal{A},Q,g)$, with the sequence of tile types $Q=(\Theta_{1},\Theta_{2},..,\Theta_{\zeta})$. A trajectory $\Psi=(A_{0},A_{1},A_{2},..)$ is said to be complete with respect to $(\mathcal{A},Q,g)$ if the following hold.
@@
      Define $G_{A}(A)$ of a configuration $A$ to be a graph whose vertices are non-empty coordinated tiles and edges are drawn between neighbouring coordinated tiles. Let a cut be defined in the usual graph-theoretic way, as a partition of graph vertices into two disjoint subsets. The set of cut-edges is then the set of edges with one vertex in each partition. Then, any cut of $G_{A}(A_{t})$ is such that the sum of interaction strengths over cut-edges $\sum_{i}g(\sigma_{i},\sigma_{i}^{\prime})>0$, where $\sigma_{i}$ and $\sigma_{i}^{\prime}$ are neighbouring faces corresponding to an edge. A configuration $A_{t}$ that obeys this assumption for some interaction function $g$ is called g-valid. Informally, a g-valid configuration is one in which any subconfiguration of tiles in a configuration will be bound to the rest of the configuration with a total strength of at least 1.
    4. (d)